# ConeGS: Error-Guided Densification Using Pixel Cones for Improved Reconstruction with Fewer Primitives

### v1

@@ 5 Evaluation
Figure 6: PSNR, LPIPS and FPS plots for a specified primitive budget (left) and without a budget (right), where the number of primitives corresponds to the chosen $\beta$ values. Numerical results are provided in the appendix. Values are averaged across Mip-NeRF360 [3] and OMMO [33].
@@
Ours
10k iters
Ours
20k iters
Ours
40k iters
EDGS [27]
3DGS [24]
(SfM init.)
MCMC [25]
(rand. init.)
MCMC [25]
(SfM init.)
@@
Table 2: Quantitative results showing overall scene reconstruction timings for 3DGS [24], MCMC [25] and EDGS [27] methods, as well as our method with differing iNGP reconstruction durations, with the number of Gaussians capped at 11M. Results are averaged over the Mip-NeRF360 [3] dataset. 3DGS time reports the target scene optimization, including densification, while init. time shows the iNGP reconstruction needed by our method and the initial matching for EDGS [27].
@@ 5.2 Ablations
We analyze the impact of individual components in our method by conducting a series of ablation studies, presented in Table 3. (a), (b) Longer iNGP reconstruction only slightly improves reconstruction quality. (c) continuing training the iNGP model also during optimization, (d) initializing Gaussians with the ground truth pixel color, or (e) predicting spherical harmonics with iNGP, leads to marginally worse results. Demonstrating the strength of our densification method, using (f) pixel-cone-sized primitives during initialization, or even not using any initialization (g), results in worse PSNR but maintains low LPIPS and considerably improves rendering speed, thanks to less overlap between primitives, reducing blending. (h) Sampling pixels uniformly instead of guiding Gaussian creation using the $L_{1}$ loss from the training set produces lower reconstruction quality, although due to uniform sampling in image space still focusing more on parts of the scene seen most across views, the drop is not drastic. Densifying with 3DGS depth (i), also without using iNGP even for initialization (j), strongly affects the results. Similarly, k-NN sizing of newly added primitives based on their closest neighbors (k), or changing the opacity penalty (l), (m), (n), has a large effect on reconstruction quality, reinforcing the benefits of our densification approach. (o), (p) Altering the size of Gaussians created during densification from their default pixel-width cone size slightly reduces reconstruction quality, although the small difference suggests that Gaussians are quickly resized to fit the scene.
@@ 6 Conclusion
We introduce ConeGS, a reconstruction pipeline that replaces cloning-based densification with a method guided by photometric error and a coarse iNGP proxy, where new primitives are sized by pixel cones. Together with an improved opacity penalty, this allows creating primitives independently of existing structures through more flexible exploration. ConeGS consistently improves reconstruction quality and rendering performance across Gaussian budgets, with strong gains under tight primitive constraints. It achieves up to a 0.6 PSNR increase and a $20\%$ speedup over cloning-based baselines.
Limitations: ConeGS works well on standard scenes but can struggle with very large scenes, inaccurate camera poses, or sparse viewpoints, sometimes producing extra floating Gaussians (see appendix). Its gains are also limited at very high Gaussian budgets, where dense coverage reduces the benefit of error-guided placement, offering mainly faster rendering.
@@
ConeGS (iNGP retrain)
29.27 0.893 0.168 212 29
@@
Tanks & Temples [26]
Truck 24.54 / 0.822 / 0.291 / 184 23.59 / 0.803 / 0.318 / 128 23.86 / 0.810 / 0.316 / 155 23.33 / 0.799 / 0.314 / 183
@@
Deep Blending [19]
Dr Johnson 28.78 / 0.875 / 0.338 / 447 23.58 / 0.808 / 0.440 / 521 28.16 / 0.869 / 0.343 / 550 27.66 / 0.865 / 0.347 / 444
@@
Tanks & Temples [26]
Truck 25.41 / 0.858 / 0.204 / 122 24.80 / 0.842 / 0.253 / 93 25.24 / 0.851 / 0.242 / 68 24.66 / 0.850 / 0.212 / 114
@@
Deep Blending [19]
Dr Johnson 29.20 / 0.890 / 0.291 / 270 28.85 / 0.881 / 0.307 / 151 28.71 / 0.884 / 0.313 / 180 28.60 / 0.888 / 0.292 / 269
@@
Tanks & Temples [26]
Truck 25.34 / 0.864 / 0.181 / 97 25.06 / 0.851 / 0.234 / 70 25.57 / 0.862 / 0.217 / 45 25.02 / 0.862 / 0.181 / 82
@@
Deep Blending [19]
Dr Johnson 28.98 / 0.886 / 0.282 / 194 28.85 / 0.884 / 0.293 / 106 29.06 / 0.884 / 0.291 / 164 28.75 / 0.890 / 0.277 / 180
@@
Tanks & Temples [26]
Truck 25.43 / 0.865 / 0.167 / 69 25.39 / 0.859 / 0.217 / 53 25.86 / 0.869 / 0.193 / 38 25.09 / 0.865 / 0.162 / 55
@@
Deep Blending [19]
Dr Johnson 29.01 / 0.884 / 0.274 / 127 28.75 / 0.886 / 0.282 / 77 28.83 / 0.882 / 0.285 / 104 28.65 / 0.887 / 0.268 / 113

### v2

@@ 5 Evaluation
Figure 6: PSNR, LPIPS and FPS plots with (left) and without (right) a primitive budget, where counts correspond to $\beta$ values. Averaged across Mip-NeRF360 [3] and OMMO [33]. Numerical results are provided in the appendix.
@@
Ours 10k iters Ours 20k iters Ours 40k iters EDGS [27] 3DGS [24] (SfM init.) MCMC [25] (rand. init.) MCMC [25] (SfM init.)
@@
Table 2: Quantitative results averaged over Mip-NeRF360 [3] showing reconstruction timings for 3DGS [24], MCMC [25], EDGS [27], and our method (with varying iNGP durations), capped at 1M Gaussians. 3DGS time reports target scene optimization and densification, while init. time shows iNGP reconstruction for our method and initial matching for EDGS [27].
@@ 5.2 Ablations
We analyze each component’s impact through ablation studies in Table 3. (a), (b) Longer iNGP reconstruction only slightly improves reconstruction quality. (c) continuing training the iNGP model also during optimization, (d) initializing Gaussians with the ground truth pixel color, or (e) predicting spherical harmonics with iNGP, leads to marginally worse results. Demonstrating the strength of our densification method, using (f) pixel-cone-sized primitives during initialization, or even not using any initialization (g), results in worse PSNR but maintains low LPIPS and considerably improves rendering speed, thanks to less overlap between primitives, reducing blending. (h) Sampling pixels uniformly instead of guiding Gaussian creation using the $L_{1}$ loss from the training set produces lower reconstruction quality, although due to uniform sampling in image space still focusing more on parts of the scene seen most across views, the drop is not drastic. Densifying with 3DGS depth (i), also without using iNGP even for initialization (j), strongly affects the results. Similarly, k-NN sizing of newly added primitives based on their closest neighbors (k), or changing the opacity penalty (l), (m), (n), has a large effect on reconstruction quality, reinforcing the benefits of our densification approach. (o), (p) Altering Gaussian sizes from their default pixel-width cone size during densification slightly reduces quality, yet the small difference suggests Gaussians quickly resize to fit the scene.
@@ 6 Conclusion
We introduce ConeGS, a reconstruction pipeline replacing cloning-based densification with a method guided by photometric error and a coarse iNGP proxy, where new primitives are sized by pixel cones. Together with an improved opacity penalty, this allows creating primitives independently of existing structures through more flexible exploration. ConeGS consistently improves reconstruction quality and rendering performance across Gaussian budgets, with strong gains under tight primitive constraints. It achieves up to 0.6 PSNR increase and 20% speedup over cloning-based baselines.
Limitations: ConeGS performs well on standard scenes but may struggle with large-scale environments, inaccurate poses, or sparse views, occasionally creating floaters (see appendix). Benefits are also reduced at high Gaussian budgets, where dense coverage limits the impact of error-guided placement, primarily offering faster rendering.
Acknowledgments
Stefano Esposito acknowledges travel support from the European Union’s Horizon 2020 research and innovation program under ELISE Grant Agreement No. 951847.
@@ Acknowledgments
ConeGS (iNGP retrain) 29.27 0.893 0.168 212 29
@@
Tanks & Temples [26] Truck 24.54 / 0.822 / 0.291 / 184 23.59 / 0.803 / 0.318 / 128 23.86 / 0.810 / 0.316 / 155 23.33 / 0.799 / 0.314 / 183
@@
Deep Blending [19] Dr Johnson 28.78 / 0.875 / 0.338 / 447 23.58 / 0.808 / 0.440 / 521 28.16 / 0.869 / 0.343 / 550 27.66 / 0.865 / 0.347 / 444
@@
Tanks & Temples [26] Truck 25.41 / 0.858 / 0.204 / 122 24.80 / 0.842 / 0.253 / 93 25.24 / 0.851 / 0.242 / 68 24.66 / 0.850 / 0.212 / 114
@@
Deep Blending [19] Dr Johnson 29.20 / 0.890 / 0.291 / 270 28.85 / 0.881 / 0.307 / 151 28.71 / 0.884 / 0.313 / 180 28.60 / 0.888 / 0.292 / 269
@@
Tanks & Temples [26] Truck 25.34 / 0.864 / 0.181 / 97 25.06 / 0.851 / 0.234 / 70 25.57 / 0.862 / 0.217 / 45 25.02 / 0.862 / 0.181 / 82
@@
Deep Blending [19] Dr Johnson 28.98 / 0.886 / 0.282 / 194 28.85 / 0.884 / 0.293 / 106 29.06 / 0.884 / 0.291 / 164 28.75 / 0.890 / 0.277 / 180
@@
Tanks & Temples [26] Truck 25.43 / 0.865 / 0.167 / 69 25.39 / 0.859 / 0.217 / 53 25.86 / 0.869 / 0.193 / 38 25.09 / 0.865 / 0.162 / 55
@@
Deep Blending [19] Dr Johnson 29.01 / 0.884 / 0.274 / 127 28.75 / 0.886 / 0.282 / 77 28.83 / 0.882 / 0.285 / 104 28.65 / 0.887 / 0.268 / 113